OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation −9 + 8v = 7 + 4v is given below. Choose the step that is missing from the solution. −9 + 8v = 7 + 4v 4v = 16 v = 4 WHICH ANSWER IS IT? : 8v = −2 + 4v −9 + 12v = 7 8v = 16 + 4v −2 + 8v = 4v

OpenStudy (anonymous):

i would go with this one 8v = 16 + 4v but the question makes no sense to me

OpenStudy (anonymous):

if you start with \[−9 + 8v = 7 + 4v\] and then add \(9\) to both sides you get \[8v=16+4v\]

OpenStudy (anonymous):

then subtract \(4v\) from both sides which gives the next step \[4v=16\]

OpenStudy (anonymous):

but this supposes that there is some right way to solve the problem, which there is not it is a really really bad question

OpenStudy (anonymous):

"missing step" is your math teachers education

OpenStudy (anonymous):

I know, I am so lost haha with this question. Especially since english is my second language. But thank you for your help.

OpenStudy (anonymous):

math is your math teachers fifth language yw